# Photonic Doping of Epsilon-Near-Zero Bragg Microcavities

### v1

@@ 3.4 Magnetic Purcell Factor Enhancement in NC-Doped ENZ Cavity
To quantify the spontaneous-emission enhancement in the NC-doped ENZ Bragg cavity, we evaluate the PF and compare it with that of a single NC or an array of NCs embedded in an infinite homogeneous medium. The formulation follows the Green-tensor approach previously introduced for two-dimensional (2D) photonic systems 37, extended here to account for both electric and magnetic dipole emitters by means of the corresponding electric and magnetic Green functions.
For an electric dipole $\mathbf{p}=p_{0}\,\mathbf{u}_{p}$, with orientation $\mathbf{u}_{p}$ at position $\mathbf{r}_{0}$, the normalized decay rate is given by the LDOS expression 38
@@
Using the relation between the Green tensor and the local electric field (Appendix C), one obtains the relation 2
@@
A parallel derivation applies to a magnetic dipole $\mathbf{m}=m_{0}\,\mathbf{u}_{m}$, with orientation $\mathbf{u}_{m}$ at $\mathbf{r}_{0}$. Starting from the magnetic LDOS and invoking the relation between $\mathbf{H}$ and the magnetic Green tensor (Appendix C), we obtain
@@
Figure 14: (a) Real and imaginary parts of the effective refractive index of TE and TM modes, together with their corresponding field distributions, in a BR cavity with $N=14$ pairs of cladding layers on each side of a half-wave core, loaded with an array of cylindrical NCs of radius $R=150~\mathrm{nm}$ and refractive index $n=3.5$, with a period of $p=750~\mathrm{nm}$. (b) Transmission spectrum of the same cavity for TE (solid curve) and TM (dashed curve) modes.
@@
Figure 14a compares the calculated PF spectra at the center of a single NC in free space (dashed line), an array of 25 NCs in free space, and the same NC array embedded in the Bragg cavity (solid line). The embedded array exhibits a nearly 35-fold Purcell enhancement relative to the isolated NC or free-space array. The corresponding electric and magnetic field distributions at resonance are shown in Figs. 14b and 14c. In these simulations, the Bragg cavity consists of $N=14$ cladding-layer pairs on each side of a half-wave core, loaded with an array of 25 NCs of radius $R=150~\mathrm{nm}$ and refractive index $n=3.5$, with period $p=750~\mathrm{nm}$. The cavity width is $W=18.75~\mu\mathrm{m}$.
The magnetic hot-spot formation and Purcell enhancement effects discussed so far for arrays of particles are not limited to circular-cylinder geometries. As illustrated in Fig. 15, replacing the circular cross-section NCs of Fig. 14 with square cross-section NCs yields similar magnetic hot spots and even stronger PF values, exceeding $10^{3}$. The side length of each square NC is $300~\mathrm{nm}$, while all other structural parameters are kept identical to those in Fig. 14.
Figure 15: Magnetic Purcell factor spectrum for an array of square cross-section NCs with side length $300~\mathrm{nm}$ embedded in the Bragg cavity, with all other structural parameters identical to Fig. 14. A pronounced resonance is observed near $1505~\mathrm{nm}$, where the PF exceeds $10^{3}$. The insets show the corresponding electric ($|E|$) and magnetic ($|H|$) field distributions at resonance, highlighting strong magnetic hot-spot confinement inside the NCs.
@@
The Purcell enhancement effect can also be observed in higher-order NZI modes of the cavity, which arise from reflections at the cavity sidewalls due to imperfect impedance matching with the surrounding medium, thereby forming standing-wave patterns inside the cavity. Interestingly, by tuning the cavity width, the PF at the center of cavity, associated with these higher-order modes can be optimized. For example, as shown in Fig. 16a, for an array of 25 NCs ($R=150~\mathrm{nm}$ and $p=500~\mathrm{nm}$) embedded in a Bragg cavity with $N=14$ and a width of $W=12.9~\mu\mathrm{m}$, the PF reaches values as high as 4600 for a higher-order NZI TE0-like mode, with the corresponding magnetic field profile shown in Fig. 16c.
@@
Figure 16: Purcell enhancement in higher-order NZI modes of a Bragg cavity with $N=14$ cladding-layer pairs and width $W=12.9~\mu\mathrm{m}$, loaded with an array of 25 NCs of radius $R=150~\mathrm{nm}$ and period $p=500~\mathrm{nm}$. (a) Magnetic Purcell factor spectrum showing multiple resonances, with a maximum PF $\sim 4600$ at $\lambda=1356.4~\mathrm{nm}$. (b) Magnetic field distribution $\Re(H_{z})$ at $\lambda=1372~\mathrm{nm}$ corresponding to a lowest-order cavity mode. (c) Magnetic field distribution $\Re(H_{z})$ at $\lambda=1356.4~\mathrm{nm}$, corresponding to a highest-order NZI TE0-like mode responsible for the giant PF enhancement.

### v2

@@ 3.4 Magnetic Purcell Factor Enhancement in NC-Doped ENZ Cavity
To quantify the spontaneous-emission enhancement in the NC-doped ENZ Bragg cavity, we evaluate the PF and compare it with that of a single NC or an array of NCs embedded in an infinite homogeneous medium. The formulation follows the Green-tensor approach previously introduced for two-dimensional (2D) photonic systems [Qiao2011], extended here to account for both electric and magnetic dipole emitters by means of the corresponding electric and magnetic Green functions.
For an electric dipole $\mathbf{p}=p_{0}\,\mathbf{u}_{p}$, with orientation $\mathbf{u}_{p}$ at position $\mathbf{r}_{0}$, the normalized decay rate is given by the LDOS expression [Novotny2012]
@@
Using the relation between the Green tensor and the local electric field (SI, Sec. C), one obtains the relation [Lobet2020]
@@
A parallel derivation applies to a magnetic dipole $\mathbf{m}=m_{0}\,\mathbf{u}_{m}$, with orientation $\mathbf{u}_{m}$ at $\mathbf{r}_{0}$. Starting from the magnetic LDOS and invoking the relation between $\mathbf{H}$ and the magnetic Green tensor (SI, Sec. C), we obtain
@@
Figure 13a compares the calculated PF spectra at the center of a single silicon NC in free space (dashed line), an array of 25 NCs in free space (dashed-dotted line), and the same NC array embedded in the Bragg cavity (solid line). The embedded array exhibits a nearly 35-fold Purcell enhancement relative to the isolated NC or free-space array. The corresponding electric and magnetic field distributions at resonance are shown in Figs. 13b and 13c. In these simulations, the Bragg cavity consists of $N=14$ cladding-layer pairs on each side of a half-wave core, loaded with an array of 25 silicon NCs of radius $R=150~\mathrm{nm}$ with period $p=750~\mathrm{nm}$. The cavity width is $W=18.75~\mu\mathrm{m}$.
Figure 13: (a) Magnetic PF for a single NC in free space, an array of 25 NCs in free space, and the same NC array embedded in the Bragg cavity. (b) Electric-field magnitude $|E|$ and (c) magnetic-field component $\mathrm{Re}(H_{z})$ at the resonant wavelength. The Bragg cavity comprises of $N=14$ cladding-layer pairs on each side of a half-wave core and is loaded with an array of 25 NCs ($R=150~\mathrm{nm}$, $n=n_{\mathrm{Si}}$, $p=750~\mathrm{nm}$).
@@
The Purcell enhancement effect can also be observed in higher-order NZI modes of the cavity, which arise from reflections at the cavity sidewalls due to imperfect impedance matching with the surrounding medium, thereby forming standing-wave patterns inside the cavity. Interestingly, by tuning the cavity width, the PF at the center of cavity, associated with these higher-order modes can be optimized. For example, as shown in Fig. 14a, for an array of 25 silicon NCs ($R=150~\mathrm{nm}$ and $p=500~\mathrm{nm}$) embedded in a Bragg cavity with $N=14$ and a width of $W=14~\mu\mathrm{m}$, the PF reaches values as high as 5600 for a higher-order NZI TE0-like mode, with the corresponding magnetic field profile shown in Fig. 14c.
@@
Figure 14: Purcell enhancement in higher-order NZI modes of a Bragg cavity with $N=14$ cladding-layer pairs and width $W=14~\mu\mathrm{m}$, loaded with an array of 25 NCs ($R=150~\mathrm{nm}$) and period $p=500~\mathrm{nm}$. (a) Magnetic Purcell factor spectrum showing multiple resonances, with a maximum PF $\approx 5600$ at $\lambda=1356.5~\mathrm{nm}$. (b) Magnetic field distribution $\Re(H_{z})$ at $\lambda=1372~\mathrm{nm}$ corresponding to a lowest-order cavity mode. (c) Magnetic field distribution $|H|$ at $\lambda=1356.5~\mathrm{nm}$, corresponding to a highest-order NZI TE0-like mode responsible for the giant PF enhancement.